OpenStudy (anonymous):

Perform the requested operation or operations. f(x) = x minus five divided by eight.; g(x) = 8x + 5, find g(f(x)). @LXelle

OpenStudy (lxelle):

\[f(x)= \frac{ (x-5) }{ 8}\] \[g(x)= 8x+5\]

OpenStudy (lxelle):

g is the first function that you wanted. so what you do is you substitute f(x) into g(x). get what i mean?

OpenStudy (lxelle):

\[g(f(x)) = 8(\frac{ x-5 }{ 8 }) + 5\] then you solve it from there.

OpenStudy (anonymous):

is it going to equal g(f(x)) = 8x + 35 @LXelle

OpenStudy (anonymous):

or i can say x-5/8

OpenStudy (lxelle):

No, no. Dont assume.

OpenStudy (anonymous):

Is it g(f(x)) = x - five divided by eight.

OpenStudy (lxelle):

No, never. How are you getting back the f(x)?

OpenStudy (anonymous):

g(f(x)) = x that might be the answer

OpenStudy (lxelle):

\[g(f(x))=8(\frac{ x-5 }{ 8 })+5\]

OpenStudy (lxelle):

From there cancel of the 8 leaving only \[x-5 + 5\]

OpenStudy (lxelle):

Therefore what you get is only x left.

OpenStudy (lxelle):

Do you know why I cancel the 8s?

OpenStudy (anonymous):

no why

OpenStudy (lxelle):

you factorize. 8 divided by 8 is?

OpenStudy (anonymous):

1

OpenStudy (lxelle):

what is left is gonna be 1(x-5)+5 Agree?

OpenStudy (lxelle):

so when you open the bracket you get x-5+5 agree?

OpenStudy (anonymous):

x-10

OpenStudy (lxelle):

nooooo. -5+5 is not 10.

OpenStudy (anonymous):

0

OpenStudy (lxelle):

Yeah so you got your answer.

OpenStudy (lxelle):

the answer is is only x.

OpenStudy (anonymous):

thnx

OpenStudy (lxelle):

:)